OpenStudy (anonymous):

y=x^+6X+8 ......I WANT THIS EQUATION IN TO FACTORED FORM !

OpenStudy (anonymous):

y=(x+2)(x+4)

OpenStudy (anonymous):

just think of it as two numbers that times to make 8 and also add to make 6 then put them in to the brackets..so 2x4=8, 2+4=6

OpenStudy (anonymous):

y=(x+3)^-1 in to standard form.. how do u do that\/?

OpenStudy (anonymous):

then y= 1(x^+3x+3x+9)-1